# Guido Franchetti1 and Derek Harland2

### v1

@@ 2.1 Preliminaries
In the particular case $2p\in\mathbb{N}$, charge $n$ hyperbolic monopoles coincide with circle-invariant instantons on $S^{4}$ with second Chern number $2np$ [1]. This follows from the fact that the metric on $S^{1}\times H^{3}$ is conformally equivalent to the metric on $S^{4}\setminus S^{2}$, with the circle action on $S^{1}\times H^{3}$ corresponding to rotations of $S^{4}$ that fix the removed $S^{2}$. Hyperbolic monopoles can be identified with circle-invariant anti-self-dual connections $A+\Phi\,d\theta$ on $S^{1}\times H^{3}$. Since the anti-self-dual equations are conformally invariant, any circle-invariant instanton on $S^{4}$ gives rise to a solution of the Bogomolny equation on $S^{1}\times H^{3}$.
The conformal equivalence of $S^{1}\times H^{3}$ and $S^{4}\setminus S^{2}$ can be realised explicitly as follows. We model $S^{4}$ as the set of unit vectors in $\mathbb{R}^{5}$. By writing $y_{1}=r_{1}\cos\theta$, $y_{2}=r_{1}\sin\theta$ with $r_{1}>0$, the metric on $\mathbb{R}^{5}\setminus\mathbb{R}^{3}$ is seen to be conformally equivalent to the metric on $S^{1}\times H^{4}$:
The equation defining $S^{4}$, $r_{1}^{2}+y_{3}^{2}+y_{4}^{2}+y_{5}^{2}=1$, determines an embedding $H^{3}\subset H^{4}$. We thus obtain a conformal identification of $S^{4}\setminus S^{2}$ with $S^{1}\times H^{3}$. The circle group acts naturally on both. The excluded set $S^{2}\subset S^{4}$, corresponding to $r_{1}=0$, is the fixed point set of the circle action and is naturally identified with the boundary of $H^{3}$.
A circle-invariant instanton $B$ on $S^{4}$ pulls back to a circle-invariant anti-self-dual connection on $S^{1}\times H^{3}$. One can find a gauge transformation $g:S^{1}\times H^{3}\to SU(2)$ such that
$g^{-1}Bg+g^{-1}dg=\Phi\,d\theta+A,$ (8)
with $(A,\Phi)$ a monopole on $H^{3}$, and $\theta$ a coordinate on $S^{1}$. Since $B$, $\Phi$ and $A$ are all invariant under $\theta\mapsto\theta+2\pi$, $g(\theta+2\pi)g(\theta)^{-1}$ is parallel with respect to $B$. Since $B$ is irreducible, this means that $g(\theta+2\pi)g(\theta)^{-1}=\pm I_{2}$. The holonomy of $B$ about $S^{1}$ is therefore given by ±$\pm\exp(-2\pi\Phi)$. Since the holonomy tends to $\mathrm{Id}$ as one approaches the fixed $S^{2}\subset S^{4}$, the eigenvalues of $\Phi$ approach ±$\pm ip$ for some $p\in\tfrac{1}{2}\mathbb{N}$. Thus circle-invariant instantons give rise to hyperbolic monopoles with half-integer mass parameter $p$; the converse to this statement was proved in [26, 27].
For later use, we now describe the circle action on the bundle $V\to S^{4}$ in more detail. Over $S^{1}\times H^{3}$, we have a trivialisation $V=S^{1}\times H^{3}\times\mathbb{C}^{2}$ and the circle action is $(\theta,x,s)\mapsto(\theta+t,x,s)$. The gauge transformation $g$ in (8) must take the form
$g=\begin{pmatrix}e^{-i\theta p}&0\\ 0&e^{i\theta p}\end{pmatrix}$ (9)
with respect to the splitting $V=L^{+}\oplus L^{-}$ over $S^{2}_{\infty}$, because $B(\frac{\partial}{\partial\theta})\to 0$ as one approaches this fixed $S^{2}$. Therefore, after making a gauge transformation with $g$, the circle action on the bundle $L^{+}\oplus L^{-}$ over $S^{1}\times S^{2}_{\infty}$ is
$(\theta,x,s_{+},s_{-})\mapsto(\theta+t,x,e^{-itp}s_{+},e^{itp}s_{-}).$ (10)
Therefore the circle group acts with weights pminus-or-$\mp p$ on the bundles $L^{\pm}$ over the $S^{2}\subset S^{4}$ that is fixed by the circle action. Strictly speaking, if $2p$ is odd the group that acts on $V\to S^{4}$ is a double cover of the circle group.
2.2 The framed and unframed moduli spaces
The Bogomolny equation is invariant under gauge transformations, and it is natural to study the moduli space of solutions modulo gauge transformations. In fact, there are two natural moduli spaces to consider, the framed and unframed moduli spaces. We define an (unframed) gauge transformation to be a function $g:H^{3}\to SU(2)$ that extends to a smooth circle-invariant section of the bundle $\operatorname{Aut}(V)\to S^{4}$. Note that circle-invariance demands that on the fixed $S^{2}\subset S^{4}$, $g$ maps $L^{+}$ to $L^{+}$ and $L^{-}$ to $L^{-}$. The moduli space of unframed monopoles is the space $M^{u}_{n,p}$ of monopoles with charge $n$ and mass parameter $p$ that extend smoothly to $S^{4}$, modulo unframed gauge transformations. This is known to be a smooth manifold of dimension $14n-$ 1 [7].
To define framed monopoles, we need to recall the definition of the Chern–Simons number. Let $A^{0},A^{1}$ be two connections on $H^{3}$ and let $\tilde{A}=(1-t)A^{0}+tA^{1}$ be a connection on $[0,1]\times H^{3}$. The Chern-Simons number of $A_{0},A_{1}$ is defined to be
$CS[A^{0},A^{1}]=2\int_{[0,1]\times H^{3}}\operatorname{Tr}(F^{\tilde{A}}\wedge F^{\tilde{A}}).$ (11)
@@ 2.2 The framed and unframed moduli spaces
Let $A^{0},A^{1},A^{2}$ be connections satisfying the boundary conditions of a monopole, and let $g:H^{3}\to SU(2)$ be a gauge transformation. Then $CS[A^{0},A^{1}]=CS[g^{-1}A^{0}g+g^{-1}dg,g^{-1}A^{1}g+g^{-1}dg]$. Moreover,
$CS[A^{0},A^{2}]=CS[A^{0},A^{1}]+CS[A^{1},A^{2}],\quad CS[A^{1},A^{0}]=-CS[A^{0},A^{1}]$ (13)
@@
Gauge invariance is immediate from the definition (11), as $F^{\tilde{A}}\mapsto g^{-1}F^{\tilde{A}}g$ under gauge transformations. The first condition of (13) can be proved from the explicit formula (12):
$\begin{split}CS[A^{0},A^{1}]+&CS[A^{1},A^{2}]-CS[A^{0},A^{2}]\\ &=\int_{S^{2}_{\infty}}\operatorname{Tr}(A^{0}\wedge A^{1}+A^{1}\wedge A^{2}-A^{0}\wedge A^{2})\\ &=\int_{S^{2}_{\infty}}\operatorname{Tr}(A^{1}\wedge A^{1}+(A^{0}-A^{1})\wedge(A^{1}-A^{2})).\end{split}$ (14)
The first of these two terms vanishes by cyclicity of the trace and skew-symmetry of the wedge product. The second vanishes because $A^{i}-A^{j}$ respects the splitting $V_{S^{2}_{\infty}}=L^{+}\oplus L^{-}$, in other words, it take values in an abelian subalgebra $\mathfrak{u}(1)\subset\mathfrak{su}(2)$. Thus the first identity of (13) holds. The second is immediate from either equation (11) or (12). ∎
We say that two connections $A^{1},A^{2}$ are framed gauge equivalent if they are gauge equivalent and $CS[A^{1},A^{2}]=0$. The moduli space of framed monopoles is the space $M^{f}_{n,p}$ of monopoles with charge $n$ and mass parameter $p$ that extend smoothly to $S^{4}$, modulo framed gauge equivalence.
There is a natural map $M^{f}_{n,p}\to M^{u}_{n,p}$ which sends a framed monopole to its (unframed) gauge equivalence class. Give a choice of reference monopole $(A^{0},\Phi^{0})$, there is also a natural map $M^{f}_{n,p}\to\mathbb{R}$ which sends $(A,\Phi)$ to $CS[A^{0},A]$. Together these give a map
$M^{f}_{n,p}\to\mathbb{R}\times M^{u}_{n,p}$ (15)
which we claim is a bijection. The map is injective because, by definition, if $(A^{1},\Phi^{1})$ and $(A^{2},\Phi^{2})$ are gauge equivalent and $CS[A^{0},A^{1}]=CS[A^{0},A^{2}]$ they are framed gauge equivalent by equation (13). To show that the map is surjective we use the following lemma, which also shows that the unframed gauge transformations are the ones generated by the Higgs field:
@@
$CS[A,A^{s}]=-4\pi nps.$ (16)
@@
We prove the identity by direct calculation from the explicit formula (12) (one can also prove this directly from the definition (11)). We find that $\frac{d}{ds}A_{s}=d^{A_{s}}\Phi$ and hence that
$\begin{split}\frac{d}{ds}CS[A,A_{s}]&=\int_{H^{3}}\operatorname{Tr}(d^{A_{s}}\Phi\wedge dA_{s}+A_{s}\wedge d(d^{A_{s}}\Phi)+2A_{s}\wedge A_{s}\wedge d^{A_{s}}\Phi)\\ &\quad+\int_{S^{2}_{\infty}}\operatorname{Tr}(A\wedge d^{A_{s}}\Phi)\\ &=\int_{H^{3}}\big{[}2\operatorname{Tr}(d^{A_{s}}\Phi\wedge F^{A_{s}})-d\operatorname{Tr}(A_{s}\wedge d^{A_{s}}\Phi)\big{]}\\ &\quad+\int_{S^{2}_{\infty}}\operatorname{Tr}(A\wedge d^{A_{s}}\Phi)\\ &=-4\pi np.\end{split}$ (17)
In going from the second to the final line we applied Stokes’ theorem, used the fact that $d^{A^{s}}\Phi|_{S^{2}_{\infty}}=0$ since the monopole abelianises at the boundary, and evaluated the remaining term using the identity (6). The result follows. ∎
To prove surjectivity of (15), let $t\in\mathbb{R}$ and let $(A,\Phi)$ be a monopole, thought of as an element of $M^{u}_{n,p}$. We need to find a gauge equivalent monopole whose Chern-Simons number is $t$. Let
$s=\frac{1}{4\pi np}(CS[A^{0},A]-t)$ (18)
Lemma 2, combined with equation (13), shows that gauge-transforming $(A,\Phi)$ with $\exp(s\Phi)$ produces a monopole $\tilde{A}$ with $CS[A^{0},\tilde{A}]=t$.
@@
Our definition of the framed moduli space is different from others in the literature, and we explain the relation with other definitions in Appendix A. Our definition is motivated by monopole dynamics. One can extend the energy functional (4) to define a lagrangian gauge theory on ×$\mathbb{R}\times H^{3}$. Dynamical configurations consist of a gauge field $A$ and scalar field $\Phi$ on ×$\mathbb{R}\times H^{3}$. Then, for fixed $t_{0},t_{1}\in\mathbb{R}$, the quantity
$\Theta(t_{0},t_{1})=2\int_{[t_{0},t_{1}]\times H^{3}}\operatorname{Tr}(F^{A}\wedge F^{A})$ (19)
@@
is clearly gauge-invariant and equal to 0 for any static configuration. By working in a gauge where $A_{t}=0$, slowly moving dynamical configurations can be approximated by paths in the monopole moduli space, parametrised by $t\in\mathbb{R}$. Consider a path of the form $(A^{t},\Phi^{t})=(g_{t}^{-1}Ag_{t}+g_{t}^{-1}dg_{t},g_{t}^{-1}\Phi g_{t})$, where $(A,\Phi)$ is a fixed monopole on $H^{3}$ and $g$ is a $t$-dependent gauge transformation. It can be shown that $\Theta(s,t)=CS[A^{s},A^{t}]$. In particular, Lemma 2 shows that, for the choice $g_{t}=\exp(t\Phi)$, $\Theta$ is non-zero, even though this path is trivial in the unframed moduli space. This example shows that in order to describe monopole dynamics it is necessary to work in the framed moduli space, rather than the unframed moduli space.
@@
Paths in the moduli space generated by gauge transformation with $\Phi$ are known as dyons and carry an electric charge, which is defined to be the abelian electric flux through $S^{2}_{\infty}$ and computed by the integral $\frac{1}{2p}\int_{S^{2}_{\infty}}\operatorname{Tr}(\Phi\ast\frac{dA}{dt})$. If $\frac{dA}{dt}=cd^{A}\Phi$ then the electric flux is $\frac{c}{2p}\int_{S^{2}_{\infty}}\operatorname{Tr}(\Phi F^{A})$ and the rate of change of the Chern-Simons number is $4c\int_{H^{3}}\operatorname{Tr}(d^{A}\Phi\wedge F^{A})$. Integration by parts then shows that the electric charge is proportional to the rate of change of the Chern-Simons number.

### v2

@@ 2.2 The framed and unframed moduli spaces
Let $A^{0},A^{1},A^{2}$ be connections satisfying the boundary conditions of a monopole, and let $g:H^{3}\to SU(2)$ be an unframed gauge transformation. The Chern–Simons number is gauge-invariant, in the sense that
$CS[A^{0},A^{1}]=CS[g^{-1}A^{0}g+g^{-1}dg,g^{-1}A^{1}g+g^{-1}dg].$ (25)
It is skew-symmetric, meaning that
$CS[A^{1},A^{0}]=-CS[A^{0},A^{1}].$ (26)
Moreover, if $A^{0},A^{1},A^{2}$ are unframed gauge equivalent then
$CS[A^{0},A^{2}]=CS[A^{0},A^{1}]+CS[A^{1},A^{2}].$ (27)
@@
Gauge invariance is immediate from the definition (23), as $F^{\tilde{A}}\mapsto g^{-1}F^{\tilde{A}}g$ under gauge transformations. Skew-symmetry is immediate from (24), cyclicity of the trace, and skew-symmetry of the wedge product. For the final result (27) we first compute
$\begin{split}CS[A^{0},A^{1}]+&CS[A^{1},A^{2}]-CS[A^{0},A^{2}]\\ &=\int_{S^{2}_{\infty}}\operatorname{Tr}(A^{0}\wedge A^{1}+A^{1}\wedge A^{2}-A^{0}\wedge A^{2})\\ &=\int_{S^{2}_{\infty}}\operatorname{Tr}((A^{1}-A^{0})\wedge(A^{2}-A^{0})-A^{0}\wedge A^{0}).\end{split}$ (28)
The second of these two terms vanishes by cyclicity of the trace and skew-symmetry of the wedge product. To analyse the first we first introduce the section $T$ of $\operatorname{End}(V)$ over $S^{2}_{\infty}$ such that $Tv=\pm v$ for $v\in L^{\pm}$. Since the connections $A^{i}$ respect the splitting $V=L^{+}\oplus L^{-}$, they make $T$ parallel. It follows that
$A^{j}=\frac{1}{2}T^{-1}dT+T\alpha^{j}$ (29)
for 1-forms $\alpha^{j}$ on $L^{+}$. By hypothesis, the connections $A^{1},A^{2}$ differ from $A^{0}$ by unframed gauge transformations $g^{j}$ with $j=1,2$. Over $S^{2}_{\infty}$, these gauge transformations take the form $g^{j}=\exp(i\lambda^{j}T)$ for functions $\lambda^{j}:S^{2}_{\infty}\to\mathbb{R}$. So
$A^{j}-A^{0}=iTd\lambda^{j}$ (30)
over $S^{2}_{\infty}$. It follows from (28) that
$CS[A^{0},A^{1}]+CS[A^{1},A^{2}]-CS[A^{0},A^{2}]=-2\int_{S^{2}_{\infty}}d\lambda^{1}\wedge d\lambda^{2}\\ =-2\int_{S^{2}_{\infty}}d(\lambda^{1}\wedge d\lambda^{2})=0.$ (31)
We say that two connections $(A^{1},\Phi^{1}),(A^{2},\Phi^{2})$ are framed gauge equivalent if they are unframed gauge equivalent and $CS[A^{1},A^{2}]=0$. It follows from the previous lemma that framed gauge equivalence is an equivalence relation. We define the moduli space $M^{f}_{n,p}$ of framed monopoles to be the quotient of $C_{n,p}$ by framed gauge equivalence.
There is a natural map
$M^{f}_{n,p}\to M^{u}_{n,p}$ (32)
that sends a framed monopole to its unframed gauge equivalence class. There is also natural action of $\mathbb{R}$ on the framed moduli space that is generated by infinitesimal gauge transformation with $\Phi$. Concretely, the action of $s\in\mathbb{R}$ is
$(A,\Phi)\mapsto(A^{s},\Phi):=(g_{s}^{-1}dg_{s}+g_{s}^{-1}Ag_{s},\Phi);\quad g_{s}:=\exp(s\Phi).$ (33)
This clearly preserves the fibres of (32), because $g_{s}$ is an unframed gauge transformation. The following lemma shows that this action is free.
@@
$\frac{dA^{s}}{ds}=d^{A^{s}}\Phi$ (34)
and
$CS[A,A^{s}]=-8\pi nps.$ (35)
In particular, $A^{s}$ is framed gauge equivalent to $A$ if and only if $s=0$.
@@
We calculate
$\frac{d}{ds}g_{s}=g_{s}\Phi,\quad\frac{d}{ds}g_{s}^{-1}=-\Phi g_{s}^{-1}\quad\text{and}\quad\frac{d}{ds}dg_{s}=g_{s}d\Phi+dg_{s}\Phi.$ (36)
So
$\frac{d}{ds}A^{s}=g_{s}^{-1}(g_{s}d\Phi+dg_{s}\Phi+Ag_{s}\Phi)-\Phi g_{s}^{-1}(dg_{s}+A_{s}\Phi)=d^{A^{s}}\Phi.$ (37)
@@
We prove (35) by direct calculation from the explicit formula (24) (one can also prove this directly from the definition (23)),
$\begin{split}\frac{d}{ds}CS[A,A^{s}]&=\int_{H^{3}}\operatorname{Tr}(d^{A^{s}}\Phi\wedge dA^{s}+A^{s}\wedge d(d^{A^{s}}\Phi)+2A^{s}\wedge A^{s}\wedge d^{A^{s}}\Phi)\\ &\quad+\int_{S^{2}_{\infty}}\operatorname{Tr}(A\wedge d^{A^{s}}\Phi)\\ &=\int_{H^{3}}\big{[}2\operatorname{Tr}(d^{A^{s}}\Phi\wedge F^{A^{s}})-d\operatorname{Tr}(A^{s}\wedge d^{A^{s}}\Phi)\big{]}+\int_{S^{2}_{\infty}}\operatorname{Tr}(A\wedge d^{A^{s}}\Phi)\\ &=\int_{H^{3}}2\operatorname{Tr}(d^{A^{s}}\Phi\wedge F^{A^{s}}).\end{split}$ (38)
since the boundary conditions imply that $d^{A^{s}}\Phi=0$ on $S^{2}_{\infty}$. Therefore, by (12),
$\frac{d}{ds}CS[A,A^{s}]=-8\pi np$ (39)
The result follows. ∎
Proposition 10.
The moduli space of framed monopoles is naturally an $\mathbb{R}$-principal bundle over the moduli space of unframed monopoles.
Proof.
We have already exhibited a projection (32) from $M^{f}_{n,p}$ to $M^{u}_{n,p}$ and an action (33) of $\mathbb{R}$ on the fibres. We showed in the previous lemma that the action is free. It remains to show that the action is transitive.
Let $(A^{0},\Phi^{0})$ and $(A^{1},\Phi^{1})$ be two monopoles that are unframed gauge equivalent. We seek an $s$ such that the action (33) of $s$ on $(A^{0},\Phi^{0})$ gives a monopole which is framed gauge equivalent to $(A^{1},\Phi^{1})$. Let $(A^{s},\Phi^{s})=(g_{s}^{-1}A^{0}g_{s}+g_{s}^{-1}dg_{s},\Phi^{0})$ denote the image of $(A^{0},\Phi^{0})$ under this action. Using lemma (9), we can choose $s$ such that
$CS[A^{0},A^{s}]=CS[A^{0},A^{1}].$ (40)
Since $A^{0},A^{1},A^{s}$ are unframed gauge equivalent, we can apply the identity (27) to show that
$CS[A^{s},A^{1}]=CS[A^{0},A^{1}]-CS[A^{0},A^{s}]=0.$ (41)
So $(A^{s},\Phi^{s})$ and $(A^{1},\Phi^{1})$ are framed gauge equivalent. ∎
Corollary 11.
$\dim(M^{f}_{n,p})=4n$.
Proof.
Proposition 10 shows that $\dim(M^{f}_{n,p})=\dim(M^{u}_{n,p})+1$, and Braam [6] has shown that shown that $\dim(M^{u}_{n,p})=4n-1$. ∎
Our definition of the framed moduli space is different from the definition given in Atiyah–Hitchin [2] for euclidean monopoles. We compare the two definitions in appendix A. Atiyah–Hitchin’s framed moduli space involves a choice of point in $S^{2}_{\infty}$, so it does not obviously admit an action of the isometry group of $H^{3}$. For euclidean monopoles, they showed that their framed moduli space does admit an action of the euclidean isometry group [2], but their argument does not adapt easily to the hyperbolic setting. In contrast, our definition leads to a framed moduli space that clearly admits an action of the isometry group of $H^{3}$.
@@
Our definition of the framed moduli space is motivated by monopole dynamics. One can extend the energy functional (10) to define a lagrangian gauge theory on $\mathbb{R}\times H^{3}$. Dynamical configurations consist of a gauge field $A$ and scalar field $\Phi$ on $\mathbb{R}\times H^{3}$. Then, for fixed $t_{0},t_{1}\in\mathbb{R}$, the quantity
$\Theta(t_{0},t_{1})=\int_{[t_{0},t_{1}]\times H^{3}}\operatorname{Tr}(F^{A}\wedge F^{A})$ (42)
@@
is clearly gauge-invariant and equal to 0 for any static configuration (i.e. any configuration that is pulled back from $H^{3}$). By working in a gauge where $A_{t}=0$, slowly moving dynamical configurations can be approximated by paths in the monopole moduli space, parametrised by $t\in\mathbb{R}$. Consider a path of the form $(A^{t},\Phi^{t})=(g_{t}^{-1}Ag_{t}+g_{t}^{-1}dg_{t},g_{t}^{-1}\Phi g_{t})$, where $(A,\Phi)$ is a fixed monopole on $H^{3}$ and $g$ is a $t$-dependent gauge transformation. It can be shown that $\Theta(s,t)=CS[A^{s},A^{t}]$. In particular, Lemma 9 shows that, for the choice $g_{t}=\exp(t\Phi)$, $\Theta$ is non-zero. So this is not a static monopole. The corresponding path $\mathbb{R}\to M^{f}_{n,p}$ is non-constant, but the corresponding path $\mathbb{R}\to M^{u}_{n,p}$ is constant. So the unframed moduli space is unable to distinguish this dynamical monopole from a static monopole, but the framed moduli space is able to do so.
@@
Paths in the moduli space generated by gauge transformations with $\Phi$ are known as dyons and carry an electric charge, which is defined to be the abelian electric flux through $S^{2}_{\infty}$ and computed by the integral $\frac{1}{2p}\int_{S^{2}_{\infty}}\operatorname{Tr}(\Phi\ast\frac{dA}{dt})$. For $(A^{t},\Phi^{t})=(g_{t}^{-1}Ag_{t}+g_{t}^{-1}dg_{t},g_{t}^{-1}\Phi g_{t})$, by (34) $\frac{dA^{t}}{dt}=d^{A^{t}}\Phi$. Denoting by $F^{t}=g_{t}^{-1}Fg_{t}$ the curvature of $A_{t}$, the electric flux is $\frac{1}{2p}\int_{S^{2}_{\infty}}\operatorname{Tr}(\Phi^{t}F^{t})=\frac{1}{2p}\int_{S^{2}_{\infty}}\operatorname{Tr}(\Phi F)=-2\pi n$ and by (35) the rate of change of the Chern-Simons number is $8\pi np$. Hence the electric charge is proportional to the rate of change of the Chern-Simons number.
For much of this article we will work with tangent vectors to the framed (and unframed) moduli spaces. The tangent spaces can be described as follows. First, a tangent vector to the configuration space $C_{n,p}$ at a point $(A,\Phi)$ consists of an $\mathfrak{su}(2)$-valued 1-form $a$ and 0-form $\phi$ solving the linearised Bogomolny equation,
$d^{A}\phi+[a,\Phi]=\ast d^{A}a,$ (43)
such that $b=a+\phi\,d\theta$ extends to a smooth circle-invariant section of $\Lambda^{1}(S^{4})\otimes\operatorname{End}(V)$. Given any $\mathfrak{su}(2)$-valued function $\chi$ that extends to a circle-invariant section of $\operatorname{End}(V)$ over $S^{4}$, $(a,\phi):=(d^{A}\chi,[\Phi,\chi])$ is a solution of (43) representing an infinitesimal unframed gauge transformation. Thus we define the space $\mathfrak{g}^{u}_{A,\Phi}$ of infinitesimal unframed gauge transformations to be the set of all pairs $(d^{A}\chi,[\Phi,\chi])$ such that $\chi:H^{3}\to\mathfrak{su}(2)$ extends smoothly to a circle-invariant section of $\operatorname{End}(V)\to S^{4}$.
In order to define infinitesimal framed gauge transformations we look at the linearisation of the framing condition. Let $g_{s}$ be a family of framed gauge transformations such that $g_{0}=1$, $g_{s}^{-1}\partial_{s}g_{s}|_{s=0}=\chi$ and set $A^{s}=g_{s}^{-1}Ag_{s}+g_{s}^{-1}dg_{s}$. Since $g_{s}$ is framed, $CS[A,A^{s}]=0$. A computation similar to the one in the proof of Lemma 9 then shows that
$0=\left.\frac{d}{ds}\right|_{s=0}CS[A,A^{s}]=\int_{H^{3}}2\operatorname{Tr}(F^{A}\wedge d^{A}\chi).$ (44)
Hence we define the space $\mathfrak{g}^{f}_{A,\Phi}$ of infinitesimal framed gauge transformations to be
$\mathfrak{g}^{f}_{A,\Phi}=\left\{(d^{A}\chi,[\Phi,\chi])\in\mathfrak{g}^{u}_{A,\Phi}:\int_{H^{3}}\operatorname{Tr}(F^{A}\wedge d^{A}\chi)=0\right\}.$ (45)
Both $\mathfrak{g}^{u}_{A,\Phi}$ and $\mathfrak{g}^{f}_{A,\Phi}$ are subspaces of the space $T_{A,\Phi}C_{n,p}$ of solutions $(a,\phi)$ of (43) that extend smoothly to $S^{4}$ in a circle-invariant way. The tangent spaces to the framed and unframed moduli spaces are the quotient spaces
$T_{A,\Phi}M^{f}_{n,p}=T_{A,\Phi}C_{n,p}/\mathfrak{g}^{f}_{A,\Phi},\quad T_{A,\Phi}M^{u}_{n,p}=T_{A,\Phi}C_{n,p}/\mathfrak{g}^{u}_{A,\Phi}.$ (46)